# Cost-aware Stopping for Bayesian Optimization

### v3

@@ Appendix B Theoretical Analysis and Calculations
In the following lemma, we prove a point-wise lower bound on the expected improvement before stopping for our recommended pairing of PBGI/LogEIPC stopping rule paired with either PBGI or LogEIPC acquisition function. [^4] In fact, any acquisition function (e.g., expected improvement-cost (EIC) (Hu et al., 2025)) that ensures the one-step expected improvement is worth the evaluation cost before the stopping time $\tau$ can apply here.
@@
We treat the two algorithms—meaning, the two acquisition function and stopping rule pairs—together and write $\mathcal{F}_{t}=\sigma\!\left(\{x_{i},y_{i}\}_{i=1}^{t}\right)$ for the filtration generated by the observations. Since we are minimizing, the one–step improvement after iteration $t$ is $y_{1:t-1}^{*}-y_{1:t}^{*}$, where recall that $y_{1:t}^{*}=\min_{1\leq i\leq t}y_{i}$. By our assumption about the random function $f$, $\mathbb{E}[y_{1:1}^{*}]=\mu$ and the quantity $y_{1:1}^{*}-\min\nolimits_{x\in X}f(x)$ has finite expectation $U<\infty$. Denote the posterior expected improvement function as $\alpha_{t}^{\mathrm{EI}}(x)=\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;y^{*}_{1:t}\right)$.
@@
$=\mu-\mathbb{E}\!\left[\min_{x\in X}f(x)\right]+\mathbb{E}[c(x_{1})]$
@@
since $\mathbb{E}[y_{1}]=\mu$, $C=c(x_{1})$, and $U=\mu-\mathbb{E}\left[\min_{x\in X}f(x)\right]$. This is exactly (11). ∎
@@
$\mathbb{E}\left[\sum_{t=1}^{\tau}\lambda\cdot c(x_{t})\right]\leq\mathbb{E}\!\left[y^{*}_{1:\tau}-\min_{x\in X}f(x)+\sum_{t=1}^{\tau}\lambda\cdot c(x_{t})\right]\leq\mu-\mathbb{E}\left[\min_{x\in X}f(x)\right]+\lambda C=U+\lambda C.$ (18)
@@
Note 4.
In this paper, we focus on the setting where $f$ is drawn from a Gaussian process, i.e., $f\sim\mathcal{GP}(\mu,K)$. In this case, the term $U=\mu-\mathbb{E}\!\left[\min_{x\in X}f(x)\right]$ can be further bounded above using classical results on the expected supremum/infimum of Gaussian processes; see, for example, Lifshits (2012, Theorem 10.1).
@@ Appendix C Experimental Setup
All experiments are implemented based on BoTorch (Balandat et al., 2020). Each Bayesian optimization procedure is initialized with $2(d+1)$ random samples, where $d$ is the dimension of the search domain. For Bayesian regret experiments, we follow the standard practice to generate the initial random samples using a quasirandom Sobol sequence. For empirical experiments, we randomly sample configuration IDs from a fixed pool of candidates—2,000 for LCBench and 32,768 for NATS-Bench. All computations are performed on CPU.
@@
For all experiments, we follow the standard practice to apply Matérn kernels with smoothness $5/2$ and length scales learned from data via maximum marginal likelihood optimization, and standardize input variables to be in $0,1]^{d}$. For empirical experiments, we standardize output variables to be zero-mean and unit-variance, but not for Bayesian regret experiments. In this work, we consider the noiseless setting and set the fixed noise level to be $10^{-6}$. In the unknown-cost experiments, we follow Astudillo et al. (2021) to model the objective and the logarithm of the cost function using independent Gaussian processes.
@@
For the 1D Bayesian regret experiments, we optimize over 10,001 grid points. For the 8D Bayesian regret experiment, we use BoTorch’s ’gen_candidates_torch’, a gradient-based optimizer for continuous acquisition function maximization, as it avoids reproducibility issues caused by internal randomness in the default scipy optimizer. For the empirical experiments, since LCBench and NATS-Bench provide only 2,000 and 32,768 configurations respectively, we optimize the acquisition function by simply applying an argmin/argmax over the acquisition values of the unevaluated configurations, without using any gradient-based methods.
@@
For PBGI, we follow Xie et al. (2024) to compute the Gittins indices using $100$ iterations of bisection search without any early stopping or other performance and reliability optimizations.
For UCB/LCB-based acquisition functions and stopping rules, we follow the original GP-UCB paper Srinivas et al. (2009) and the choice in UCB/LCB based stopping rules (Makarova et al., 2022; Ishibashi et al., 2023) to use the schedule $\beta_{t}=2\log(dt^{2}\pi^{2}/6\delta)$, where $d$ is the dimension. We also adopt their choice of $\delta=10^{-1}$ and a scale-down factor of $5$.
@@
For PRB, we follow Wilson (2024) to use the schedule $N_{t}=\max(\lceil 64*1.5^{t-1}\rceil,1000)$ for number of posterior samples, risk tolerance $\delta=0.05$. The error bound $\epsilon$ is set to be 0.1 for Bayesian regret experiments and 0.5% of the best test error (here, the misclassification rate) among all configurations for the empirical experiments.
For the 8D Bayesian regret experiments, we apply moving average over 20 iterations to mitigate oscillations in the optimizer. Figure˜5 illustrates the challenges these oscillations pose when computing stopping rule statistics and shows the improvement with moving average. For consistency, we also apply 20-iteration averaging to the GSS and Convergence baselines.
Figure 5: Comparison of the raw and moving-averaged PBGI/LogEIPC stopping rule statistics (i.e., the LogEIPC acquisition values) in Bayesian regret 8D experiments for multiple acquisition functions, with linear cost and stopping thresholds at $\log(0.1)$, $\log(0.01)$ and $\log(0.001)$. Left: The unaveraged statistic exhibits large, high-frequency fluctuations due to the difficulty of acquisition optimization in high dimensions. Right: Applying moving average (window=20) smooths these wiggles, yielding a more stable stopping signal.
@@
We omit the KG acquisition function and stopping rule due to its high computational cost, as they are shown to be computationally intensive in the runtime experiments of Xie et al. (2024).
@@
In the periodic cost setting, the evaluation cost fluctuates across the domain. Following Astudillo et al. (2021), we define the periodic cost as
@@
A visualization of the three cost functions is provided in Figure˜6.
@@
In the known-cost experiments, for LCBench, we use 0.001 times the number of model parameters as a proxy for runtime. This proxy is motivated by our observation of an approximately linear relationship between the number of model parameters and the actual runtime, with slope close to 0.001 (see Figure 7). Importantly, the number of model parameters can be computed in advance, before the Bayesian optimization procedure, based on the network structure and classification task, as we explain in detail below.
@@ D.2 Order of Stopping and Posterior Updates
Following the discussions in Section˜3, we always compute our proposed stopping rule with respect to the optimal acquisition function value of the next round—namely, the one which is obtained after posterior updates have been performed. One could alternatively consider checking the stopping criteria before posterior updates, which is backward-looking rather than forward-looking. Figure˜10 provides an empirical comparison between the two choices, showing that stopping after the posterior update leads to stronger empirical performance.
This suggests that the theoretical guarantee for the Gittins index policy in the correlated Pandora’s Box setting by Gergatsouli & Tzamos (2023), which is based on the before-posterior-update stopping, could potentially be improved by adopting the after-posterior-update stopping.
@@ D.3 Additional Experiment Results: Empirical
This subsection presents additional experiment results on hyperparameter optimization over the three LCBench datasets and neural architecture size search over the three NATS datasets.
To isolate the effect of the acquisition function on cost-adjusted regret, we report the simple regret of several acquisition functions in the budgeted setting. We compare LogEIPC, PBGI, LCB, and TS, and additionally include PBGI-D with our recommended choice $\lambda_{0}=U/(B-C)$ from Section˜3.1.1, where $U=50$ (reflecting the [0,100] range of classification accuracy) and BC=10, [^0] B-C=10{,}000 (corresponding to a budget after initial evaluation of 10,000 seconds, or roughly 3 hours). Cost-aware methods (LogEIPC and the PBGI variants) outperform cost-unaware ones (UCB and TS), with PBGI at smaller $\lambda$ consistently better than LogEIPC. This mirrors the findings of Xie et al. (2024) and helps explain the strong performance of the matched PBGI combination in our main experiments.
@@
We count the number of trials in which a stopping rule fails to trigger within our iteration cap of 200 and present the results in Table˜1. From the table, we observe that on datasets from the NATS benchmark, regret-based and acquisition-based stopping rules—except for ours—often fail to stop early. On LCBench datasets, some regret-based stopping rules such as SRGap-med and UCB–LCB also frequently exceed the cap. In contrast, our stopping rule consistently stops early, which aligns with our theoretical result in Corollary˜3.
@@
As shown in the empirical setting in Figures˜12, 13, 14, 15, 16 and 17, cost-adjusted regret generally decreases in the early iterations and then increases. The turning point is exactly the hindsight-optimal fixed-iteration stopping point, and our PBGI/LogEIPC stopping rule consistently performs close to this optimum, particularly when paired with the PBGI acquisition function.
@@
Figure 18: Comparison of cost-adjusted simple regret on LCBench with $\lambda=10^{-4}$, using scaled proxy runtime vs. scaled actual runtime as evaluation cost. While our PBGI/LogEIPC stopping rule performs slightly worse under actual runtime (e.g., dropping from best to second-best on the higgs dataset), likely due to cost model mismatch, it remains close to the hindsight optimal in all cases.
@@
Astudillo et al. (2021, Proposition 2) proposed modeling unknown cost $c(x)$ via
@@
The difference in sign before the variance term reflects how each formulation handles predictive uncertainty: (22) encourages more exploration than (23). For PBGI under the unknown-cost setting, it is more natural to replace $c(x)$ in (5) with $\mathbb{E}[c(x)]$ using (23), as this aligns with how costs enter the root-finding problem. For LogEIPC, both variants are possible—we refer to the (22) version as LogEIPC-inv and the (23) version as LogEIPC-exp. However, equivalence between PBGI and LogEIPC stopping rules and our theoretical guarantees hold only with (23) but not with (22). Accordingly, we use (23) for both methods in our experiments to maintain consistency and preserve this equivalence. Figure˜19 shows performance of acquisition function and stopping rule pairs under the unknown-cost setting, which are qualitatively similar to the known-cost setting.
Figure 19: Comparison of cost-adjusted simple regret across acquisition function and stopping rule pairs under the unknown-cost setting on LCBench with $\lambda=10^{-4}$. Our PBGI/LogEIPC stopping rule remains close to the hindsight optimal when paired with the LogEIPC-exp or PBGI acquisition function, sometimes slightly worse than heuristics such as GSS and Convergence. It is also slightly worse than Immediate on Adult, likely due to a cost-model mismatch in the unknown-cost setting.

### v4

@@ Appendix B Theoretical Analysis and Calculations
In the following lemma, we prove a point-wise lower bound on the expected improvement before stopping for our recommended pairing of PBGI/LogEIPC stopping rule paired with either PBGI or LogEIPC acquisition function. [^5] In fact, any acquisition function (e.g., expected improvement-cost (EIC) (hu2025adjusted)) that ensures the one-step expected improvement is worth the evaluation cost before the stopping time $\tau$ can apply here.
@@
We treat the two algorithms—meaning, the two acquisition function and stopping rule pairs—together and write $\mathcal{F}_{t}=\sigma\!\left(\{x_{i},y_{i}\}_{i=1}^{t}\right)$ for the filtration generated by the observations. Since we are minimizing, the one–step improvement after iteration $t$ is $y_{1:t-1}^{*}-y_{1:t}^{*}$, where recall that $y_{1:t}^{*}=\min_{1\leq i\leq t}y_{i}$. By our assumption about the random function $f$, $\mathbb{E}[y_{1:1}^{*}]=\mu(x_{1})$ and the quantity $y_{1:1}^{*}-\min\nolimits_{x\in X}f(x)$ has finite expectation $U<\infty$. Denote the posterior expected improvement function as $\alpha_{t}^{\mathrm{EI}}(x)=\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;y^{*}_{1:t}\right)$.
@@
$=\mu(x_{1})-\mathbb{E}\!\left[\min_{x\in X}f(x)\right]+\mathbb{E}[c(x_{1})]$
@@
since $\mathbb{E}[y_{1}]=\mu(x_{1})$, $C=c(x_{1})$, and $U=\mu(x_{1})-\mathbb{E}\left[\min_{x\in X}f(x)\right]$. This is exactly (11). ∎
Remark 5.
The quantities $U$ and $C$ depend on the choice of the initial point $x_{1}$. Choosing $x_{1}$ to minimize $\mu(x_{1})+c(x_{1})$ yields a tighter bound.
Remark 6.
In practice, one may center the objective by replacing $f(x)$ with $f(x)-\mu(x)$ and then using a zero-mean GP prior. This changes only the parameterization of the posterior mean (which can be shifted back by $\mu(x)$) and does not affect the implementation of the acquisition functions or stopping rules.
@@
Immediately by Theorem˜2 and the fact that $y^{*}_{1:\tau}-\min_{x\in X}f(x)>0$ pointwise, we have that the expected cumulative cost of the algorithm
$\mathbb{E}\left[\sum_{t=1}^{\tau}c(x_{t})\right]\leq U+C.$ (18)
By Markov’s inequality, for any $\delta\in(0,1)$,
$\Pr\left[\sum_{t=1}^{\tau}c(x_{t})\leq\frac{U+C}{\delta}\right]\geq 1-\delta.$
Since $c(x_{t})\geq c_{0}$ for all $t<\tau$, it follows that
$\tau=\sum_{t=1}^{\tau}1\leq\sum_{t=1}^{\tau}\frac{c(x_{t})}{c_{0}}=\frac{1}{c_{0}}\sum_{t=1}^{\tau}c(x_{t}).$
Therefore,
$\Pr\left[\tau\leq\frac{U+C}{\delta\cdot c_{0}}\right]\geq\Pr\left[\frac{1}{c_{0}}\sum_{t=1}^{\tau}c(x_{t})\leq\frac{U+C}{\delta\cdot c_{0}}\right]=\Pr\left[\sum_{t=1}^{\tau}c(x_{t})\leq\frac{U+C}{\delta}\right]\geq 1-\delta.$
See 4
Proof.
@@
$\mathbb{E}\left[\sum_{t=1}^{\tau}\lambda\cdot c(x_{t})\right]\leq\mathbb{E}\!\left[y^{*}_{1:\tau}-\min_{x\in X}f(x)+\sum_{t=1}^{\tau}\lambda\cdot c(x_{t})\right]\leq\mu(x_{1})-\mathbb{E}\left[\min_{x\in X}f(x)\right]+\lambda C=U+\lambda C.$ (19)
@@
Remark 7.
In this paper, we focus on the setting where $f$ is drawn from a Gaussian process, i.e., $f\sim\mathcal{GP}(\mu(\cdot),K)$. In this case, the term $U=\mu(x_{1})-\mathbb{E}\!\left[\min_{x\in X}f(x)\right]$ can be further bounded above using classical results on the expected supremum/infimum of Gaussian processes; see, for example, lifshits2012lectures.
@@ Appendix C Experimental Setup
All experiments are implemented based on BoTorch (balandat2020botorch). Each Bayesian optimization procedure is initialized with $2(d+1)$ random samples, where $d$ is the dimension of the search domain. For Bayesian regret experiments, we follow the standard practice to generate the initial random samples using a quasirandom Sobol sequence. For empirical experiments, we randomly sample configuration IDs from a fixed pool of candidates—2,000 for LCBench and 32,768 for NATS-Bench. All computations are performed on CPU.
@@
For all experiments, we follow the standard practice to apply Matérn kernels with smoothness $5/2$ and length scales learned from data via maximum marginal likelihood optimization, and standardize input variables to be in $0,1]^{d}$. For empirical experiments, we standardize output variables to be zero-mean and unit-variance, but not for Bayesian regret experiments. In this work, we consider the noiseless setting and set the fixed noise level to be $10^{-6}$. In the unknown-cost experiments, we follow astudillo2021multi to model the objective and the logarithm of the cost function using independent Gaussian processes.
@@
For the 1D Bayesian regret experiments, we optimize over 10,001 grid points. For the 8D Bayesian regret experiment, we use BoTorch’s ‘gen_candidates_torch’, a gradient-based optimizer for continuous acquisition function maximization, as it avoids reproducibility issues caused by internal randomness in the default scipy optimizer. For the empirical experiments, since LCBench and NATS-Bench provide only 2,000 and 32,768 configurations respectively, we optimize the acquisition function by simply applying an argmin/argmax over the acquisition values of the unevaluated configurations, without using any gradient-based methods.
@@
For PBGI, we follow xie2024cost to compute the Gittins indices using $100$ iterations of bisection search without any early stopping or other performance and reliability optimizations.
For UCB/LCB-based acquisition functions and stopping rules, we follow the original GP-UCB paper srinivas2009gaussian and the choice in UCB/LCB based stopping rules (makarova2022automatic; ishibashi2023stopping) to use the schedule $\beta_{t}=2\log(dt^{2}\pi^{2}/6\delta)$, where $d$ is the dimension. We also adopt their choice of $\delta=10^{-1}$ and a scale-down factor of $5$.
@@
For PRB, we follow wilson2024stopping to use the schedule $N_{t}=\max(\lceil 64*1.5^{t-1}\rceil,1000)$ for number of posterior samples, risk tolerance $\delta=0.05$. The error bound $\epsilon$ is set to be 0.1 for Bayesian regret experiments and 0.5% of the best test error (here, the misclassification rate) among all configurations for the empirical experiments.
For the 8D Bayesian regret experiments, for all acquisition-value-based stopping rules, we apply moving average over 20 iterations to mitigate the fluctuations due to the imperfect acquisition function optimization. Figure˜4 illustrates the challenges these oscillations pose when computing stopping rule statistics and shows the improvement with moving average. For consistency, we also apply the 20-iteration averaging to the non-acquisition-value-based GSS and Convergence baselines.
Figure 4: Comparison of the raw and moving-averaged PBGI/LogEIPC stopping rule signals (i.e., the LogEIPC acquisition values) in Bayesian regret 8D experiments for multiple acquisition functions, with linear cost and stopping thresholds at $\log(0.1)$, $\log(0.01)$ and $\log(0.001)$. Left: The unaveraged signals exhibits large, high-frequency fluctuations due to the difficulty of acquisition optimization in high dimensions. Right: Applying moving average (window=20) smooths these wiggles, yielding more stable stopping signals.
@@
We omit the KG acquisition function and stopping rule due to its high computational cost, as they are shown to be computationally intensive in the runtime experiments of xie2024cost.
@@
In the periodic cost setting, the evaluation cost fluctuates across the domain. Following astudillo2021multi, we define the periodic cost as
@@
A visualization of the three cost functions is provided in Figure˜5.
@@
In the known-cost experiments, for LCBench, we estimate the runtime cost from the number of model parameters $p$. Specifically, for the first three datasets in the six-dataset lite version, we use 0.001 times the number of model parameters as a proxy for runtime. This proxy is motivated by our observation of an approximately linear relationship between the number of model parameters and the actual runtime, with slope close to 0.001 (see Figure 6). For the full 35-dataset version, where the slope varies across datasets, we instead use a regression-derived coefficient $\alpha$ and the proxy cost is $\alpha p$. Importantly, the number of model parameters can be computed in advance, before the Bayesian optimization procedure, based on the network structure and classification task, as we explain in detail below.
@@ D.2 Order of Stopping and Posterior Updates
Following the discussions in Section˜3, we always compute our proposed stopping rule with respect to the optimal acquisition function value of the next round—namely, the one which is obtained after posterior updates have been performed. One could alternatively consider checking the stopping criteria before posterior updates, which is backward-looking rather than forward-looking. Figure˜9 provides an empirical comparison between the two choices, showing that stopping after the posterior update leads to stronger empirical performance.
This suggests that the theoretical guarantee for the Gittins index policy in the correlated Pandora’s Box setting by gergatsouli2023weitzman, which is based on the before-posterior-update stopping, could potentially be improved by adopting the after-posterior-update stopping.
@@ D.3 Additional Experiment Results: Empirical
This subsection presents additional results for hyperparameter optimization on the LCBench datasets and neural architecture–size search on the three NATS-Bench datasets. For LCBench, we provide results for the first three datasets from the six-dataset lite version of the benchmark: Fashion-MNIST, adult, and higgs under alternative settings (fixed budget, actual runtime, and unknown cost), along with an alternative visualization. We also include per-dataset results under the proxy-runtime setting for the full set of 35 datasets.
Simple regret under the fixed-budget setting.
To isolate the effect of the acquisition function on cost-adjusted regret, we report the simple regret of several acquisition functions in the fixed-budget setting. We compare LogEIPC, PBGI, LCB, and TS, and additionally include PBGI-D with our recommended choice $\lambda_{0}=U/(B-C)$ from Section˜3.1.1, where $U=50$ (reflecting the [0,100] range of classification accuracy) and BC=10, [^0] B-C=10{,}000 (corresponding to a budget after initial evaluation of 10,000 seconds, or roughly 3 hours). Cost-aware methods (LogEIPC and the PBGI variants) outperform cost-unaware ones (UCB and TS), with PBGI at smaller $\lambda$ consistently better than LogEIPC. This mirrors the findings of xie2024cost and helps explain the strong performance of the matched PBGI combination in our main experiments.
@@
We count the number of trials in which a stopping rule fails to trigger within our iteration cap of 200 and present the results in Table˜1. From the table, we observe that on datasets from the NATS benchmark, regret-based and acquisition-based stopping rules—except for ours—often fail to stop early. On LCBench datasets, some regret-based stopping rules such as SRGap-med and UCB–LCB also frequently exceed the cap. In contrast, our stopping rule consistently stops early, which aligns with our theoretical result in Corollary˜4.
@@
As shown in the empirical setting in Figures˜11, 12, 13, 14, 15 and 16, cost-adjusted regret generally decreases in the early iterations and then increases. The turning point is exactly the hindsight-optimal fixed-iteration stopping point, and our PBGI/LogEIPC stopping rule consistently performs close to this optimum, particularly when paired with the PBGI acquisition function.
@@
Figure 17: Comparison of cost-adjusted simple regret on three LCBench datasets with $\lambda=10^{-4}$, using scaled proxy runtime vs. scaled actual runtime as evaluation cost. While our PBGI/LogEIPC stopping rule performs slightly worse under actual runtime (e.g., dropping from best to second-best on the higgs dataset), likely due to cost model mismatch, it remains close to the hindsight optimal in all cases.
@@
astudillo2021multi proposed modeling unknown cost $c(x)$ via
@@
The difference in sign before the variance term reflects how each formulation handles predictive uncertainty: (23) encourages more exploration than (24). For PBGI under the unknown-cost setting, it is more natural to replace $c(x)$ in (5) with $\mathbb{E}[c(x)]$ using (24), as this aligns with how costs enter the root-finding problem. For LogEIPC, both variants are possible—we refer to the (23) version as LogEIPC-inv and the (24) version as LogEIPC-exp. However, equivalence between PBGI and LogEIPC stopping rules and our theoretical guarantees hold only with (24) but not with (23). Accordingly, we use (24) for both methods in our experiments to maintain consistency and preserve this equivalence. Figure˜18 shows performance of acquisition function and stopping rule pairs under the unknown-cost setting, which are qualitatively similar to the known-cost setting.
Figure 18: Cost-adjusted simple regret across acquisition function and stopping rule pairs under the unknown-cost setting on LCBench with $\lambda=10^{-4}$. Our PBGI/LogEIPC stopping rule remains close to the hindsight optimal when paired with the LogEIPC-exp or PBGI acquisition function, sometimes slightly worse than heuristics such as GSS and Convergence. It is also slightly worse than Immediate on Adult, likely due to a cost-model mismatch in the unknown-cost setting.
Cost-adjusted simple regret of all LCBench datasets.
Due to space constraints, for LCBench, Figure˜3 in the main text reports min–max normalized cost-adjusted simple regret, defined as
$\frac{r-r_{\min}}{r_{\max}-r_{\min}},$
where $r$ denotes cost-adjusted regret of a given acquisition function–stopping rule pair, and $r_{\min}$ and $r_{\max}$ are the minimum (including the hindsight optimal) and maximum cost-adjusted regret across all pairs. This normalization enables aggregation across datasets with different regret scales.
Here we present the unnormalized bar-plot results for each LCBench dataset (OpenML dataset size in parentheses) under all three cost-scaling parameters in Figures˜19, 20 and 21. As discussed in the main text, our PBGI/LogEIPC stopping rule—when paired with either PBGI or LogEIPC—achieves competitive performance on roughly 75% of the 35 datasets. However, it performs consistently poorly across all three $\lambda$ values on the following datasets (see subplots with light grey background in Figures˜19, 20 and 21): Amazon_employee_access (32769), Australian (690), KDDCup09_appetency (50000), cnae-9 (1080), credit-g (1000), fabert (8237), and vehicle (846). Three additional datasets, jasmine (2984), mfeat-factors (2000) and sylvine (5124), show acceptable performance only when $\lambda=10^{-5}$.
With the exception of Amazon_employee_access (32769) and KDDCup09_appetency (50000), all datasets on which our recommended pairs (matched PBGI and matched LogEIPC) perform poorly have fewer than 10,000 instances. This suggests that relatively small dataset size may contribute to model misspecification and degraded performance.
Figure 19: Cost-adjusted simple regret of all 35 LCBench datasets when $\lambda=10^{-3}$. Datasets where matched PBGI and matched LogEIPC underperform are highlighted in light grey.
Figure 20: Cost-adjusted simple regret of all 35 LCBench datasets when $\lambda=10^{-4}$. Datasets where matched PBGI and matched LogEIPC underperform are highlighted in light grey.
Figure 21: Cost-adjusted simple regret of all 35 LCBench datasets when $\lambda=10^{-5}$. Datasets where matched PBGI and matched LogEIPC underperform are highlighted in light grey.